OpenStudy (mitu12):

simplify the following radicals square root of 96

OpenStudy (misssunshinexxoxo):

\[\sqrt{96} = \sqrt{16 } \times \sqrt6 = 4\sqrt6\]

OpenStudy (mitu12):

where do you get the 16

OpenStudy (misssunshinexxoxo):

Oops I meant \[\sqrt96 = \sqrt16\times 6\]

OpenStudy (misssunshinexxoxo):

By simplifying

OpenStudy (misssunshinexxoxo):

2 * 2 * 2 * 2 * 2 * 3 = 96

OpenStudy (mitu12):

how would i know what to number will multiply to give me 96

OpenStudy (mitu12):

i meant two

OpenStudy (misssunshinexxoxo):

Do you know how to factor?

OpenStudy (mitu12):

yes

OpenStudy (misssunshinexxoxo):

What's the square root of 16?

OpenStudy (misssunshinexxoxo):

96 divided by 6?

OpenStudy (mitu12):

4

OpenStudy (misssunshinexxoxo):

96 divided by 6 also?

OpenStudy (mitu12):

16

OpenStudy (misssunshinexxoxo):

There you go^

OpenStudy (misssunshinexxoxo):

That's where it comes from

OpenStudy (mitu12):

sorry but i dont really understand it

OpenStudy (misssunshinexxoxo):

96 /6=16 times 6

OpenStudy (mitu12):

why would i divide by 6

OpenStudy (misssunshinexxoxo):

If you know how to factor shouldn't be an issue; take it step by step

OpenStudy (mitu12):

how do you just factor 96

OpenStudy (misssunshinexxoxo):

Here are the small factors of 96 1, 2, 3, 4, 6, 8

OpenStudy (mitu12):

ok

OpenStudy (mitu12):

so you basically just check and guess

OpenStudy (misssunshinexxoxo):

No, you see what best fits

OpenStudy (misssunshinexxoxo):

Are you taking Algebra?

OpenStudy (mitu12):

im starting geometry

OpenStudy (misssunshinexxoxo):

Ok make sure to go over factoring again

OpenStudy (mitu12):

alright

OpenStudy (mitu12):

thank you ill watch it

OpenStudy (misssunshinexxoxo):

The answer to this question is 4√6

OpenStudy (misssunshinexxoxo):

It's a bit tricky but be confident. Just tackle it. My pleasure to help! :)